eric is two years older than his wife canoy. in 15 years, the sum of their ages will be 1 more than thrice eric is age now. how old are they now?

Answer by ankor@dixie-net.com(22740)

let e = Eric's age
let c = Canoy's age
Write an equation for each statement, simplify as much as possible.
:
eric is two years older than his wife canoy.
e = c + 2
in 15 years, the sum of their ages will be 1 more than thrice eric is age now.
(e+15) + (c+15) = 3e + 1
e + c + 30 = 3e + 1
c + 30 - 1 = 3e - e
c + 29 = 2e
Replace e with (c+2)
c + 29 = 2(c+2)
c + 29 = 2c + 4
29 - 4 = 2c - c
25 = c
Canoy is 25 yrs old. You can find Eric's age, check your solutions in the 2nd statement